Calculate the standard free energy change for the reaction involving the decomposition of Hydrogen peroxide to water at 25°c given that the standard free energies H²O² H²O and O² are—120.4,-288.6 and 0.0kjmol–¹ respectively
1
Expert's answer
2019-11-04T06:24:02-0500

Solution.

We write down the formulas for changing the Gibbs free energy of a chemical reaction:

"\Delta G = \Delta G(products) - \Delta G(start \ material)"

Write down the equation of the chemical reaction and equalize:

"2H2O2 = 2H2O + O2"

"\Delta G = (2 \times (-288.6) + 0) - (2 \times (-120.4)) = -336.4 \ \frac{kJ}{mole}"

"\Delta G<0,"

therefore, the reaction is spontaneous. However, this reaction under normal conditions practically does not occur or with a low reaction rate. In this reaction, catalysts are used, for example, metal ions (iron, chromium), iodine or manganese oxide (4).

Answer:

"\Delta G = -336.4 \ \frac{kJ}{mole}"